The unit of ∇×H is
1
Expert's answer
2019-04-19T10:35:56-0400

The unit of H is


"\frac{A}{m}"

The unit of ∇ is


"\frac{1}{m}"

Thus, the unit of ∇×H is


"\frac{A}{m} \frac{1}{m}= \frac{A}{m^2}"